ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 34 Ronto, Andrei, Institute of Mathematic, National Academy of Sciences, Ukraine On periodic solutions of a class of differential functional equations For a class of systems of differential functional equations of the form

x′ = F x,

some theorems are given extending previous results by A. M. Samoilenko. The use is made of the “sign” conditions on the operator F, similar to that suggested earlier by J. Hale with regard to the method of L. Cesari.